OpenStudy (anonymous):

it takes 5 weeks for mice to go from a population of 100 to a population of 1000. What is the doubling time of this population?

OpenStudy (anonymous):

this is what I did, but I don't think I'm right... 1000=100e^(k5), where k is the rate of growth 10=e^(k5) ln(10)=ln(e^k5) ln(10)=k5 k=ln(10)/5 which equals 0.46 which doesn't make sense?

OpenStudy (dumbcow):

you did it correctly, but "k" is not the doubling time, its just a constant you need to solve 2 = e^(kt) 2 = e^.46t t = ln(2)/.46

OpenStudy (anonymous):

why did you put 2=e^kt? where did the 2 come from

OpenStudy (dumbcow):

"double" means 2 times as much start with 100 , you want 200 200 = 100e^kt 2 = e^kt

OpenStudy (anonymous):

ohhh! ok thanks!